# Reinforcement learning to learn quantum states for Heisenberg scaling accuracy

### v1

@@ IV Discussion & Conclusion
The single-shot measurement scheme used in this work is proposed by the single-shot measurement learning (SSML) [25, 26]. Notably, the SSML addresses following problems of the standard approach: (i) The way to reconstruct a learned state is not known. (ii) The learned state can have negative eigenvalues [5]. (iii) The scaling of infidelity is limited to ${\cal O}(n^{-3/4})$ for $n$ copies of states without additional information about quantum state [73]. Our method also does not suffer from these problems. The trained HEA immediately provides a way to reconstruct the estimated state, and the reconstructed state always satisfies the conditions of quantum state as the state reconstruction is given by the quantum circuit. Also, the average infidelity nearly achieves the scaling of statistical limit [45]. While the SSML shows that the optimal scaling in the learning up to $6$-dimensional states, our results show the optimal scaling for higher dimensional states. Implementing the SSML on a quantum computer is difficult for high-dimensional states as it uses a parameterized special unitary gate and compiling the parameterized circuit typically incurs a high computational cost. In our method, we use the Hardware efficient ansatz of which structure can be compatible with quantum computers of low connectivity, e.g., one-dimensional array of qubits allowing neighborhood interaction.
@@
A strategy of action repetition has been used in a reinforcement learning scheme for games [78]. In this context, the motivation of action repetition is to imitate human behaviors in playing games. On the other hand, we use the action repetition strategy to control the depth of decision process with the curriculum so that the agent can experience the problems of various difficulties. Reducing the depth and breadth of the decision process is a key issue in successfully applying reinforcement learning to large-scale problems [79]. (Note that, in our problem, the discretization of action space plays a role in reducing the breadth of decision process.) The efficacy of curriculum learning has been shown in meta-learning schemes for gradient descents [39] and the quantum optimization problems [36]. In the setting of our problem, the curriculum of the previous works can be realized by gradually increasing the maximum time steps of evolution strategy from a small value to a large value. However, we do not apply this curriculum to our method as the control of maximum time step directly affects the estimation accuracy of observation-value function. We estimate the value function based on the empirical cumulative reward, and we require the learning quantum state to be completed within a maximum time step to accurately compute the cumulative reward. If the maximum time step is set to a small value, the learning process will not be completed, and this hinders the calculation of cumulative reward. The inaccurate estimation of value function can bias the agent to only solve easy problem instances, the quantum states which can be learned in a short time step.
@@
For the task of learning quantum states, we propose the meta-learning scheme based on the hybrid model of quantum and classical neural networks. We perform the learning Haar-random quantum states by training the Hardware efficient ansatz (HEA) of low connectivity with the evolution strategy. To improve the learning quantum states, we use the reinforcement learning (RL), and the RL agent learns the policy to dynamically adjust the hyperparameters of evolution strategy. The trained RL agent enables the infidelity of reconstructed states to approach the scaling of Heisenberg limit [45]. Our meta-learning model demonstrates the generalizability by showing that the RL agent can be applicable to learning quantum states of higher dimension than the dimension of quantum states used for training. The generalization is possible since the structure of RL agent is independent of the dimension of quantum system. To obtain these results, we introduce the action repetition strategy (ARS) which can facilitate the training of RL agent. We expect that our meta-learning scheme including the ARS can be leveraged to improve known classical and quantum algorithms for learning quantum states, quantum control [80], quantum machine learning [81], and quantum optimization [34, 82]. Furthermore, hyperparameter optimization and improving blackbox optimization algorithms through meta-learning have broad applications across disciplines [36, 83, 84, 85, 86].
@@ Acknowledgements.
Authors thank Minyoung Lee and Taehee Lee for their support. J.J. thanks Sang Min Lee and Jeongho Bang for discussions.
Figure A1: The results of learning a thousand of random pure states obtained without the RL agent. The target success count is set to ${\cal C}_{\text{target}}=10^{4}$. We express the average of total success count and infidelity on a logarithmic scale. The total time steps are represented by ratio $t_{H}/t_{\text{max}}$, where $t_{\text{max}}=10^{4}$, $10^{5}$, and $10^{6}$ for $1$-, $2$-, and $3$-qubit state, respectively. The shaded cells are where the evolution strategy fails to learn quantum states within the maximum time step. The columns (a)–(c) are the results of the learning $1$-, $2$-, and $3$-qubit states with the $16$ actions, respectively. The column (d) is the result of the learning $2$-qubit states with the $169$ actions. The baseline values and corresponding action $a_{\text{base}}$ are represented by the red fonts.
@@ Appendix B Elements of POMDP
    ${\cal R}$ is the space of reward $r_{t}$. The reward $r_{t}=0$ if the learning quantum states is finished at time step $t$. Otherwise, $r_{t}=-1$.
@@
    $\Omega$ is the space of observation $o_{t}$. To observe the state, we use a measurement of binary outcome; success and fail. We perform the measurement until the fail outcome appears. The observation is the number of consecutive success outcomes before a fail outcome appears.
@@ Appendix C Actor-Critic algorithm
where ${t}_{\text{max}}$ is the maximum time step of ES which is set to $3\times 10^{3}$, $10^{4}$, and $2\times 10^{4}$ for $1$-, $2$-, and $3$-qubit, respectively. We introduce ${t}_{\text{max}}$ in the denominator to normalize the empirical cumulative reward. The loss function of critic network is given by the difference between the output of critic network and the estimator of observation-value function, i.e.,
@@
where $A(o_{t}):=Q_{\phi_{1}}(o_{t})-\tilde{Q}(o_{t})$ is called advantage. The parameters of actor network $\phi_{2}$ is updated by a gradient ascent. In the typical Advantage Actor Critic (A2222C) method [90], the value function and advantage are estimated by using the time difference (TD) error. In our case, we estimate these quantities based on the empirical value function obtained by a policy-guided tree search, similar to RL using Monte-Carlo tree search [79, 91]. We use the ADAM optimizer for the gradient methods [92]. The learning rate of ADAM optimizer is set to $10^{-4}$ for the learning $1$-qubit states, and $3\times 10^{-5}$ for the learning $2$- and $3$-qubit states.
@@ Appendix D Learning quantum states without RL agent
In Fig. A1, we investigate the performance of learning a thousand random pure states without using the RL agent. We choose the performance given by an action $a_{\text{base}}$ which gives the lowest total success count as the baseline. For $1$-qubit, $a_{\text{base}}=(\sigma=0.1,\eta=0.01)$. In case of $2$-qubit, $a_{\text{base}}=(\sigma=0.1,\eta=0.001)$ for the $16$ actions, and $a_{\text{base}}=(\sigma=10^{-1},\eta=10^{-2.75})$ for the $169$ actions. For $3$-qubit, $a_{\text{base}}=(\sigma=0.033,\eta=0.0001)$. The respective baseline values are represented by the red fonts in Fig. A1, and the dashed lines in Fig. 2. The shaded cells are where the evolution strategy fails to learn quantum states within the maximum time step $t_{\text{max}}$. These results imply that, without the RL agent, finding an action which can successfully complete the learning quantum states requires much amount of computational resources. For example, without the trained RL agent, learning a thousand $3$-qubit states requires more than a million time steps for most of the actions considered.

### v2

@@ IV Discussion & Conclusion
The single-shot measurement scheme used in this work is proposed by the single-shot measurement learning (SSML) [25, 26]. Notably, the SSML addresses following problems of the standard approach: (i) The way to reconstruct a learned state is not known. (ii) The learned state can have negative eigenvalues [5]. (iii) The scaling of infidelity is limited to ${\cal O}(n^{-3/4})$ for $n$ copies of states [75]. Our method also addresses these limitations: The trained HEA immediately provides a way to reconstruct the estimated state, ensuring that the reconstructed state always satisfies the conditions of a valid quantum state, as it is represented by a quantum circuit. Also, the average infidelity nearly achieves the scaling of statistical limit [47]. While the SSML shows that the optimal scaling in the learning up to $6$-dimensional states [26], our results show the optimal scaling for higher dimensional states. Implementing the SSML on a quantum computer is challenging for high-dimensional states as it uses a parameterized special unitary gate and compiling the parameterized circuit typically incurs a high computational cost. In constrast, our approach employs the hardware efficient ansatzs of which structure can be compatible with quantum computers of low connectivity such as a one-dimensional qubit array with nearest-neighbor interactions.
@@
A strategy of action repetition has been used in a reinforcement learning scheme for games [81]. In this context, the motivation behind action repetition is to imitate human behaviors in playing games. In contrast, we use the action repetition strategy to control the depth of the decision process through the curriculum so that the agent can experience the problems of various difficulties. Reducing the depth and breadth of the decision process is a key issue in successfully applying reinforcement learning to large-scale problems [82]. (Note that in our problem, discretizing the action space reduces the breadth of the decision process.) The effectiveness of curriculum learning has been shown in meta-learning schemes for gradient descent [39] and the quantum optimization problems [36]. In our setting, the curriculum of the previous works can be realized by gradually increasing the maximum time steps of evolution strategy from a small value to a large value. However, we do not apply this curriculum to our method as controlling the maximum time step directly impacts on the estimation of the observation-value function. If the maximum time step is set to a small value, the learning process may not complete, leading to inaccurate cumulative reward estimation. This inaccuracy can bias the agent to solve only easier problem instances, which are the quantum states that can be learned within a short time step.
@@
For the task of learning quantum states, we propose the meta-learning scheme based on the hybrid model of quantum and classical neural networks. Our approach trains a hardware-efficient ansatz (HEA) with low connectivity using an evolution strategy to learn Haar-random quantum states. To enhance the learning of quantum states, we use the reinforcement learning (RL), where an RL agent optimizes the hyperparameters of the evolution strategy. The trained RL agent enables the reconstructed states to achieve infidelity scaling close to the Heisenberg limit [47]. Our meta-learning model demonstrates generalizability, as the RL agent can be applied to learning quantum states of higher dimensions than those used for training. The generalization is possible since the structure of RL agent is independent of the dimension of quantum system. To obtain these results, we introduce the action repetition strategy (ARS), which facilitates RL training. We expect that our meta-learning scheme including ARS can be leveraged to enhance existing classical and quantum algorithms for quantum state learning, quantum control [83], quantum machine learning [84], and quantum optimization [34, 85]. Moreover, meta-learning for hyperparameter optimization of blackbox optimization algorithms has broad applications across multiple disciplines [36, 86, 87, 88, 89].
@@ Acknowledgements.
Authors thank Minyoung Lee and Taehee Lee for their support. J.J. thanks Sang Min Lee, Jeongho Bang, and Akash Kundu for discussions.
Figure A1: Results of learning a thousand of random pure states obtained without the RL agent. The target success count is set to ${\cal C}_{\text{target}}=10^{4}$. We express the average of total success count and infidelity as logarithmic values in base $10$. The total time steps are represented by ratio $t_{H}/t_{\text{max}}$, where $t_{\text{max}}=10^{4}$, $10^{5}$, and $10^{6}$ for $1$-, $2$-, and $3$-qubit state, respectively. The shaded cells are where the evolution strategy fails to learn quantum states within the maximum time step. The tables in the columns (a)–(c) are the results obtained by learning of $1$-, $2$-, and $3$-qubit states with the $16$ actions, respectively. The column (d) is the results of learning $2$-qubit states with the $169$ actions. The red fonts represent the baseline values and respective actions $a_{\text{base}}$.
@@ Appendix B Elements of POMDP
    ${\cal R}$ is the space of reward $r_{t}$. The reward $r_{t}=0$ if the quantum state learning is finished at time step $t$. Otherwise, $r_{t}=-1$.
@@
    $\Omega$ is the space of observation $o_{t}$. To observe the state, we use a measurement of binary outcome; success and fail. We perform the measurement until a fail outcome appears. The observation is defined as the number of consecutive success outcomes before the fail outcome.
@@ Appendix C Actor-Critic algorithm
where ${t}_{\text{max}}$ is the maximum time step of ES which is set to $3\times 10^{3}$, $10^{4}$, and $2\times 10^{4}$ for $1$-, $2$-, and $3$-qubit, respectively. We introduce ${t}_{\text{max}}$ in the denominator to normalize the empirical cumulative reward, and its value is empirically chosen. The loss function of critic network is given by the difference between the output of critic network and the estimator of observation-value function, i.e.,
@@
where $A(o_{t}):=Q_{\phi_{1}}(o_{t})-\tilde{Q}(o_{t})$ is called advantage. The parameters of actor network $\phi_{2}$ is updated by a gradient ascent.
In the typical algorithm of Advantage Actor-Critic (A2222C) [93], the value function and advantage are estimated by using the time difference (TD) error. In our case, we estimate these quantities based on the empirical value function obtained by a policy-guided tree search, similar to RL using Monte-Carlo tree search [82, 94]. We use the ADAM optimizer for the gradient methods [95]. The learning rate of ADAM optimizer is set to $10^{-4}$ for learning $1$-qubit states, and $3\times 10^{-5}$ for learning $2$- and $3$-qubit states.
@@ Appendix D Learning quantum states without RL agent
In Fig. A1, we investigate the performance of learning a thousand random pure states without using the RL agent. We choose the performance given by an action $a_{\text{base}}$ which gives the lowest total success count as the baseline. For $1$-qubit, $a_{\text{base}}=(\sigma=0.1,\eta=0.01)$. In case of $2$-qubit, $a_{\text{base}}=(\sigma=0.1,\eta=0.001)$ for the $16$ actions, and $a_{\text{base}}=(\sigma=10^{-1},\eta=10^{-2.75})$ for the $169$ actions. For $3$-qubit, $a_{\text{base}}=(\sigma=0.033,\eta=0.0001)$. The respective baseline values are represented by the red fonts in Fig. A1, and the dashed lines in Fig. 2. The shaded cells are where the evolution strategy fails to learn quantum states within the maximum time step $t_{\text{max}}$. These results imply that, without the RL agent, finding an action which can successfully complete the quantum state learning requires much amount of computational resources. For example, without the trained RL agent, learning a thousand $3$-qubit states requires more than a million time steps for most of the actions considered.
Figure A2: Results of learning $1$- and $2$-qubit mixed states. The average infidelity $\langle\bar{f}\rangle$ and the average total success count $\langle{\cal C}_{\text{total}}\rangle$ are represented on a logarithmic scale in base $10$. $\mu$ represents the amount of noise in the mixed states. The target success counts ${\cal C}_{\text{target}}$ are set to values from $10^{1}$ to $10^{5}$. The horizontal dashed lines represent the lower bound of infidelity $\bar{f}_{l}=\mu/2$. The diagonal dashed lines indicate the scaling of Heisenberg limit.
Appendix E Learning mixed states
We investigate the performance of our meta-learning scheme for a hundred random mixed states $\hat{\varrho}$. A random $N$-qubit mixed state is given by
$\hat{\varrho}=(1-\mu)|\psi\rangle\langle\psi|+\mu\frac{I}{2^{N}},$ (A7)
where $\left|\psi\right>$ is a random pure state, and $\mu$ quantifies how much noise is in the state $\hat{\varrho}$. If $\mu=0$, the state is pure, and, if $\mu=1$, the state is the maximally mixed state. Fig. A2 shows the performance of our meta-learning scheme for the random mixed states of $\mu=0$, $\mu=10^{-3}$ and $\mu=10^{-2}$. The noise degrades the performance of our scheme. The lower bound of infidelity is given by $\bar{f}_{l}=1-\langle\psi|\hat{\varrho}|\psi\rangle=\mu/2$. Our method nearly achieves the lower bound.
Appendix F Quantum state tomography
We briefly review the process of quantum state tomography (QST). Let $\hat{\varrho}_{\text{true}}$ be an unknown state to be estimated with the QST, and let $\hat{\varrho}(\vec{r})$ be a parametric model of the unknown state. To estimate $\hat{\varrho}_{\text{true}}$, measurements $M$s are performed on the multiple copies of unknown state, and the frequency of each outcome $i$ is recorded as $f_{i}$. The QST concerns how much the probabilities of model $p_{i}(\vec{r})=\mbox{Tr}\hat{\varrho}(\vec{r})\hat{M}_{i}$ differ from the frequencies. The difference can be quantified by the Kullback–Leibler (KL) divergence, $\sum_{i}f_{i}\log(f_{i}/p_{i}(\vec{r}))$. Maximum likelihood estimation finds the parameters of the minimum KL divergence, $\vec{r}_{\text{ML}}$, by maximizing the log-likelihood function ${\cal L}(\vec{r}):=\sum_{i}f_{i}\log p_{i}(\vec{r})$. In other words, the optimal parameters are obtained by
$\vec{r}_{\text{ML}}=\arg\!\max_{\vec{r}}{\cal L}(\vec{r}).$ (A8)
Thus, in QST, the state estimation problem amounts to the optimization problem of the log-likelihood function.
For the optimization, we employ a heuristic algorithm, so-called $R\varrho R$ method [96], which updates the model $\hat{\varrho}(\vec{r})$ by following rule
$\hat{\varrho}(\vec{r}_{t+1})\leftarrow\alpha\hat{\varrho}(\vec{r}_{t})+(1-\alpha)\frac{\hat{R}_{t}\hat{\varrho}(\vec{r}_{t})\hat{R}_{t}}{\mbox{Tr}\left(\hat{R}_{t}\hat{\varrho}(\vec{r}_{t})\hat{R}_{t}\right)},$ (A9)
where $\hat{R}_{t}:=\frac{1}{m}\sum_{i}(f_{i}/p_{i}(\vec{r}_{t}))\hat{\Pi}_{i}$, and $m$ is the number of measurement bases used for the estimation. The learning rate is set to $\alpha=0.5$. This method is based on the expectation maximization algorithm, and the fact that the optimal state $\hat{\varrho}(\vec{r}_{\text{ML}})$ is an eigenstate of the operator $R$.